# Bi-LAT: Bilateral Control-Based Imitation Learning via Natural Language and Action Chunking with Transformers

### v1

@@ II-A Bilateral Control-Based Imitation Learning
Many studies on bilateral control-based imitation learning employ neural network architectures such as Long Short-Term Memory (LSTM). For example, previous studies have demonstrated tasks like drawing straight lines along a ruler [10], wiping a whiteboard with an eraser [11], and thinly slicing cucumbers [12], all of which require precise force control for successful execution. More recently, Transformers have gained traction in robotic control. Notably, Bi-ACT [13] integrates Action Chunking with Transformers (ACT) [14] into bilateral control-based imitation learning. Bi-ACT predicts not only the next action but also sequences of actions, thereby addressing the issue of cumulative errors, a common challenge in imitation learning. Moreover, Bi-ACT extends bilateral control-based imitation learning by incorporating visual data, which enables flexible adaptation to objects with varying weights and shapes. Recent developments related to Bi-ACT include ALPHA-$\alpha$ [9], a low-cost bimanual robotic platform, and the data augmentation method for bilateral control-based imitation learning with images(DABI) [15], which enhances training data by downsampling both image and robot data to improve learning efficiency.
@@ II-B Imitation Learning Using Language Models
Figure 2: Bi-LAT Data Collection
@@ III-E Inference
Fig. 5 illustrates the operational diagram for Bi-LAT inference. During inference, the Bi-LAT model receives the most recent joint data from the follower robot along with corresponding camera images and a natural language instruction specifying the desired action. Based on this input, the model predicts the subsequent action chunk of leader robot’s joints information. Bi-LAT outputs action data that include joint angles, velocities, and torques for each joint. These outputs are then converted into the requisite current for each joint by calculations performed by the bilateral control system.
@@ IV-D Experiment Result
Figure 11: Data Collection of Sponge-Twisting Task using ALPHA-$\alpha$
@@
In the conventional Bi-ACT model, which was trained without language inputs, data from both ”softly” and ”strongly” demonstrations were merged into a single training set. As a result, the model treated both force modalities as equivalent aspects of the same task. As shown in Fig. 10(a), when no language instruction was provided, the inferred joint angle and torque profiles were biased toward strong force characteristics, with joint angles distributed around 2.75 rad and torques between 0.10 and 0.20 Nm. This result indicates that the Bi-ACT model is unable to differentiate between the “softly” and “strongly” actions, thereby failing to encode the nuanced force requirements for varying task conditions. These findings underscore the necessity of incorporating natural language cues to guide force modulation, ultimately enabling a more precise and context-aware imitation learning process.
@@ V Bimanual Experiment
Figure 13: Results of Sponge-Twisting Task using ALPHA-$\alpha$
@@ V-A Hardware
As shown in Fig. 12, the bimanual robots ALPHA-$\alpha$[9] were used for experiments in bimanual robotic manipulation. A total of four robots were employed, including two leader robots operated by the human operator and two follower robots. Each robot features 6 degrees of freedom (DOF) for versatile movement, and an additional DOF for the gripper, utilizing a total of 7 motors for its operation. The bilateral control cycle was set to 1000 Hz for precise movement and data collection of joint angle, velocity, and torque. Furthermore, four RGB cameras were positioned overhead, on the side, and on the right and left gripper joints of the follower robots to capture visual observations.
@@ VI CONCLUSIONS
Experiments in both unimanual and bimanual settings confirmed that Bi-LAT reliably reproduces a range of force intensities, particularly when using SigLIP as the text encoder. This precision in force modulation, integrated with language-based commands and haptic feedback, demonstrates the framework’s robustness in adapting to dynamic manipulation tasks. Our results underscore the potential of combining bilateral control-based imitation learning with natural language cues to enhance the adaptability and intuitiveness of robotic manipulation. This hybrid approach not only refines the technical capabilities of robots in handling complex objects but also aligns with the broader vision of seamless human-robot collaboration in diverse application domains such as service robotics, industrial automation, and assistive technologies.
Future work will address more diverse object properties, refine real-time language interactions, and explore large-scale datasets for enhanced generalization.

### v2

@@ II-A Bilateral Control-Based Imitation Learning
[…]
@@ II-B Imitation Learning Using Language Models
Figure 2: Bi-LAT: Data Collection
@@ III-E Inference
Fig. 5 illustrates the operational diagram for Bi-LAT inference. During inference, the Bi-LAT model receives the most recent joint data from the follower robot along with corresponding camera images and natural language instruction specifying the desired action. Based on this input, the model predicts the subsequent action chunk of leader robot’s joints information. Bi-LAT outputs action data that include joint angles, velocities, and torques for each joint. These outputs are then converted into the requisite current for each joint by calculations performed by the bilateral control system.
@@ IV-D Experiment Result
[…]
@@
In the conventional Bi-ACT model, which was trained without language inputs, data from both “softly” and “strongly” demonstrations were merged into a single training set. As a result, the model treated both force modalities as equivalent aspects of the same task. As shown in Fig. 10(a), when no language instruction was provided, the inferred joint angle and torque profiles were biased toward strong force characteristics, with joint angles distributed around 2.75 rad and torques between 0.10 and 0.20 Nm. This result indicates that the Bi-ACT model is unable to differentiate between the “softly” and “strongly” actions, thereby failing to encode the nuanced force requirements for varying task conditions. These findings underscore the necessity of incorporating natural language cues to guide force modulation, ultimately enabling a more precise and context-aware imitation learning process.
@@ V Bimanual Experiment
[…]
@@ V-A Hardware
[…]
@@ VI CONCLUSIONS
Experiments in both unimanual and bimanual settings confirmed that Bi-LAT reliably reproduces a range of force intensities, particularly when using SigLIP as the text encoder. This precision in force modulation, integrated with language-based commands and haptic feedback, demonstrates the framework’s robustness in adapting to dynamic manipulation tasks. Our results underscore the potential of combining bilateral control-based imitation learning with natural language cues to enhance the adaptability and intuitiveness of robotic manipulation. The hybrid approach enhances complex object manipulation and promotes seamless human-robot collaboration across service, industrial, and assistive robotics.
Future work will address more diverse object properties including rigid and deformable materials, incorporate varying task difficulties, refine real-time language interactions, and explore large-scale datasets for enhanced generalization.